OpenStudy (anonymous):

Solve limit using Hospital rule. lim x-> -infinity ( (2^x-2^(-x)) / 2^x+2^(-x) ) The answer from WolphramAlpha is -1, but how?

OpenStudy (anonymous):

I got this so far: \[\frac{ 2^{x}\ln(2)+2^{-x}\ln(2) }{ 2^{x}\ln(2)-2^{-x}\ln(2) } \] How do I proceed?

OpenStudy (tkhunny):

I'm wondering why we need l'Hopital. Perhaps pondering the limit as \(x \rightarrow \infty\) might contribute to understanding.

OpenStudy (anonymous):

I think it's because it's currently in the form of infinity/infinity, which is an indeterminate form.

OpenStudy (dls):

Can we write it like this? \[\LARGE \frac{4^x-1}{4^x+1}\]

OpenStudy (dls):

x->infinity

OpenStudy (dls):

@continuume

OpenStudy (anonymous):

How did you get that?

OpenStudy (dls):

\[\LARGE \lim_{x \rightarrow \infty } \frac{2^x-2^{-x}}{2^{-x}+2^x}\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait, to negative infinity

OpenStudy (dls):

yes,typo^^

OpenStudy (dls):

Try to simplify this by yourself once,just simplify

OpenStudy (anonymous):

I did, which is what I posted earlier up. But then I don't know how to proceed with that..

OpenStudy (dls):

Okay

OpenStudy (dls):

\[\LARGE \lim_{x \rightarrow \infty} \frac{2^{x}-\frac{1}{2^x}}{\frac{1}{2^{x}}+2^{x}}\]

OpenStudy (dls):

can u do it now?

OpenStudy (dls):

\[\LARGE 2^{-x}=\frac{1}{2^{x}}\] u know this right?

OpenStudy (anonymous):

Yes, I can now see how your answer came about.

OpenStudy (dls):

Glad,lets proceed.

OpenStudy (dls):

\[\LARGE \lim_{x \rightarrow \infty} \frac{4^{x}-1}{4^{x}+1}\]

OpenStudy (dls):

thats -infty sorrry dam!

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

I see the answer now!! :)

OpenStudy (dls):

\[\LARGE \lim_{x \rightarrow -\infty} \frac{4^{x}}{1+4^{x}} -\lim_{x \rightarrow -\infty} \frac{1}{1+4^{x}}\] can i write like this?

OpenStudy (anonymous):

Yes

OpenStudy (dls):

u got the answer?

OpenStudy (anonymous):

I don't know if I did it write, but I did 4 to the power of infinity, which yields 0. I don't know if that's right.

OpenStudy (anonymous):

Simplifying, I get -1.

OpenStudy (dls):

okay it must be correct then :P

OpenStudy (anonymous):

\[\frac{ 4^{\infty}-1 }{ 4^{\infty}+1 } = -1\] Not sure if correct ... :p

OpenStudy (anonymous):

I get -1 from wolphramAlpha, hmm...

OpenStudy (dls):

:S

OpenStudy (dls):

lets move ahead?

OpenStudy (anonymous):

Actually you're right,....it's indeterminate :/

OpenStudy (dls):

:)

OpenStudy (dls):

You should know,limit of a constant is constant.. and limit of a quotient is the quotient too!

OpenStudy (dls):

\[\LARGE \frac{1}{\frac{1}{{\lim_{x \rightarrow -\infty}4^{x}}} +1} - \frac{1}{{\lim_{x \rightarrow -\infty}1+4^{x}}} \]

OpenStudy (dls):

now use continuity

OpenStudy (dls):

why dont u apply L hospital after the 2nd step :P

OpenStudy (dls):

o.o sensei's here!! :D hey @zepdrix

OpenStudy (anonymous):

he's gone :(

OpenStudy (dls):

nvm

OpenStudy (dls):

sensei hai, imasu o.o

OpenStudy (anonymous):

konichiwa. I'm still a bit lost..

zepdrix (zepdrix):

\[\huge \text{:O}\]

OpenStudy (dls):

WHY U NO FOLLOW WHAT I SAY

OpenStudy (anonymous):

(╯°□°)╯︵ ┻━┻)

zepdrix (zepdrix):

Woah woah woah woah! Simmer down now! ┬─┬ ノ( ゜-゜ノ)